OpenStudy (anonymous):

Quadrilateral EFGH has coordinates E(a, 2a), F(3a, a), G(2a, 0), and H(0, 0). Find the midpoint of HG. (2a, 0) (a, 2a) (a, a) (a, 0)

OpenStudy (anonymous):

@conner143

OpenStudy (anonymous):

ok ill try

OpenStudy (bibby):

midpoint = \((\dfrac{x_1+x_2}{2},\dfrac{y_1+y_2}{2})\)

OpenStudy (anonymous):

what does a=

OpenStudy (bibby):

it doesn't have to equal anything. it's a variable. if it helps you can set it to some arbitrary number

OpenStudy (anonymous):

a is just in place of a variable :)

OpenStudy (anonymous):

oh ok

OpenStudy (bibby):

and so the midpoint of HG would be where H=(x1,y1) and G is (x2,y2) then you can plug it in

OpenStudy (anonymous):

do you want the answer or help finding it out?

OpenStudy (anonymous):

Both would be great! :)

OpenStudy (anonymous):

ok so the answer is (a,0)

OpenStudy (anonymous):

i got it bc you use the midpoint formula. so you do this 2a-0 0-0 ----- , ---- 2 2 and when you get the answers you put them together like coordinates so 2a-0 is 2a and 2a/2 would be a so x=a and you would do 0-0 which is 0 so 0/2 is o so y=0 so therefor the answer is (a,0) does that make since @alohakaika

OpenStudy (anonymous):

Yes! :) Thanks @conner143

OpenStudy (anonymous):

ok no problem anytime just ask